OpenStudy (anonymous):

3 + a = y is the same as y = 3 + a. This is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (jlastino):

Symmetric Property

OpenStudy (anonymous):

commutative property is it not?

OpenStudy (jlastino):

@redshift no...because the commutative property states that the order doesn't matter a+b=c b+a=c 3+a = y a+3 = y The symmetric property states that if a=b then b=a thus, 3+a=y y=3+a

OpenStudy (anonymous):

Thanks